What is the pH of a solution containing 0.3 M Tris(hydroxymethyl)amino-methane (free base) and 0.2 M Tris hydrochloride? pka = 8.1
Answer #1

Tris(hydroxymethyl)amino-methane is weak base and Tris hydrochloride is salt.

this is buffer solution.

for buffer solution,

Henderson equation is

POH = PKb + log [salt] / [base]

where C is the concentration of weak base.

PKa = 8.1

PKa + PKb = 14

PKb = 14 - PKa = 14 - 8.1 = 5.9

PKb = 5.9

POH = PKb + log [salt] / [base]

POH = 5.9 + log (0.2 / 0.3)

POH = 5.72

PH + POH = 14

PH = 14 - 5.72

PH = 8.28

PH of solution is 8.28.